OpenStudy (jennychan12):

Find F'(x) if F(x) = integral from 1 to 3x^2 of 1/t dt.

OpenStudy (jennychan12):

F(x) = \[\int\limits_{1}^{3x^2} \frac{ 1 }{ t } dt\]

OpenStudy (anonymous):

Fundamental theorem of calculus?

OpenStudy (jennychan12):

I know the fundemental theorum of calculus applies. But I don't know how to use it here. I would just say 1/3x^2 - 1 but I'm pretty sure that's wrong. ( I never understood how to do this.)

OpenStudy (anonymous):

Set \(y = 3x^2\) and remember that: \[ \frac{dF}{dx} = \frac{dF}{dy}\frac{dy}{dx} \]

OpenStudy (jennychan12):

wait. i think i got it. is it F'(x) = 12/x ?

OpenStudy (jennychan12):

sorry for no work. :( is it just take the derivative and plug it in?

OpenStudy (anonymous):

First of all, what is \(dy/dx\)?

OpenStudy (jennychan12):

6x

OpenStudy (anonymous):

What's \(dF/dy\)?

OpenStudy (jennychan12):

AHHHH sheet. I typed it wrong...again. It's supposed to be F(x) = \[\int\limits_{1}^{3x^12} \frac{ 1 }{ t }dt\]

OpenStudy (jennychan12):

isn't dF/dy just 1/t ? i'll stick with my old question

OpenStudy (anonymous):

It's \(1/y = 1/(3x^2)\)

OpenStudy (jennychan12):

wait what? i thought it was derviatvie of F with respect to y...?

OpenStudy (jennychan12):

oh yeah sorry i see what u did.

OpenStudy (anonymous):

\[ \frac{d}{dy}\int^y_af(t)dt = f(y)=f(y(x)) \]

OpenStudy (anonymous):

Forgetting everything, let's just do the integral? \[\huge \int\limits\limits_{1}^{3x^{12}} \frac{ 1 }{ t }dt = \ln(3x^{12})-\ln(1)=\ln(3x^{12})\] And then get its derivative?

OpenStudy (jennychan12):

oh ok i get it now... so it's just (36x^11) / ln (3x^12) then ?

OpenStudy (anonymous):

Why is there still a logarithm?

OpenStudy (jennychan12):

oh wait whoops. forgot that. my computer's slow so i thought i had already erased that.

OpenStudy (jennychan12):

yeah it's 36x^11/3x^12 which simplifes to 12/x right?

OpenStudy (anonymous):

It should be \[\huge \frac{36x^{11}}{3x^{12}}=\frac{12}{x}\] Or, cleverly, \[\large \ln(3x^{12})=\ln(3) + \ln(x^{12})\] And then get the derivative Either way, you're right :)

OpenStudy (jennychan12):

oh okay thanks. i kept not understanding that up to now.

OpenStudy (anonymous):

:)